OpenStudy (anonymous):

I'm studying for the ASVAB and fractions are the scariest thing to my existence (I'm in my mid-20's and it's been 7 years). Can you help me set this up? I don't need the answer, just the instructions :) Two dozen shuttlewingspans and four badminton rackets are to be purchased for a playground. The shuttlewingspans are priced at $.35 each and rackets at $2.75 each. The playground receives a discount of 30% from these prices. Total cost is...? I know to multiply each cost by the amount the school is buying. Where do I go from there?

OpenStudy (anonymous):

find the total cost C. discount=c*30/100 net cost=C-DISCOUNT.

OpenStudy (anonymous):

So the shuttle things are $8.40 for the two dozen and $11.00 for the rackets. So it'd be 19.20(30)/100?

zepdrix (zepdrix):

You might be able to skip a step in there if you just think about it a sec :) If the discount is 30%, how much do you end up paying? You end paying 70% of the full price right? We can write 70% as a decimal .70 So our final price that we pay is:\[\Large 19.40\times .70\]

OpenStudy (anonymous):

So I took 19.40x.70 and somehow got 135.80 but I took out both zeros so moving the decimal over one actually gives me the answer of 13.58 which is one of the answers to choose from. Thank you very much zepdrix :)

zepdrix (zepdrix):

Woops you must've multiplied by 7.0! careful! :D